# An LLM-based Readability Measurement for Unit Tests’ Context-aware Inputs

### v1

@@ Abstract
Automated test techniques usually generate unit tests with higher code coverage than manual tests. However, the readability of automated tests is crucial for code comprehension and maintenance. The readability of unit tests involves many aspects. In this paper, we focus on test inputs. The central limitation of existing studies on input readability is that they focus on test codes alone without taking the tested source codes into consideration, making them either ignore different source codes’ different readability requirements or require manual efforts to write readable inputs. However, we observe that the source codes specify the contexts that test inputs must satisfy. Based on such observation, we introduce the Context Consistency Criterion (a.k.a, C3), which is a readability measurement tool that leverages Large Language Models to extract primitive-type (including string-type) parameters’ readability contexts from the source codes and checks whether test inputs are consistent with those contexts. We have also proposed EvoSuiteC3. It leverages C3’s extracted contexts to help EvoSuite generate readable test inputs. We have evaluated C3’s performance on $409$ JAVA classes and compared manual and automated tests’ readability under C3 measurement. The results are two-fold. First, The Precision, Recall, and F1-Score of C3’s mined readability contexts are $84.4\%$, $83\%$, and $83.7\%$, respectively. Second, under C3’s measurement, the string-type input readability scores of EvoSuiteC3, ChatUniTest (an LLM-based test generation tool), manual tests, and two traditional tools (EvoSuite and Randoop) are $90\%$, $83\%$, $68\%$, $8\%$, and $8\%$, showing the traditional tools’ inability in generating readable string-type inputs. We have conducted a survey based on the questionnaires collected from 30 programmers with varied backgrounds. The results reveal that when C3 identifies readable differences between tests, programmers tend to give similar opinions of the test’s readability of C3.
@@ 1 Introduction
Figure 1: Overview of test inputs from manual and automated tests for $30$ JAVA classes (see Sec.  5.2): Manual, EvoSuiteC3 (see Sec.  4), and ChatUniTest [1] tests contain many real-world entities, such as person names, emails, URLs, and locations; EvoSuite [2] and Randoop [3] tests contain many random inputs, as well as JAVA class names, JSON strings, and UUIDs [4] since these two tools tend to use pre-defined seeds and static or dynamic information of classes under test as test inputs
Unit testing is a popular testing technique among developers. A unit test checks whether a method inside the source code behaves functionally correctly. Existing unit test generation tools (e.g., search-based tool EvoSuite  [2, 5] and feedback-directed random tool Randoop [3]) can generate unit tests that are close to and even outperform the manually written ones in many coverage criteria [6], such as branch coverage.
@@
The readability of unit tests involves many aspects, including but not limited to test method names [12], comments [13, 14], documents [15], identifiers, and assertions [16, 17]. In this paper, we focus on test inputs. Many studies [10, 18, 19, 20] and our survey (see Sec. 5.3) show that test inputs are crucial in readability. However, the random inputs generated by existing tools (EvoSuite and Randoop) cost understanding time. A former survey [10] showed that compared to strings occurring in natural languages (e.g., "mail"), random strings (e.g., "Qu5ua") require people at most $1.6$ times more comprehension time.
@@ 16 User user7 = new User("hi!", "", "");
Code 1: A constructor from Airsonic [21] with three tests differing in inputs
To deal with this issue, Afshan et al. [10] used a statistical model to predict the likelihood of the test input occurring in the real world. They formulated the likelihood as a secondary objective to guide test generation. The main issue with this approach is that its model ignores the source code when judging the input readability. However, different source codes have different requirements for input readability. Deljouyi et al. [22] proposed to extract understandable inputs from end-to-end (E2E) tests’ execution trace logs since an E2E test aims to simulate a real user to execute an application’s entire workflow [23, 24] and may contain understandable inputs. However, this approach is semi-automated since it requires human effort to write E2E tests.
@@ 1.1 A Motivating Example
We leverage an example to demonstrate the lack of readability of automated tests and the challenges of existing studies. This example is a simplified constructor method under test (MUT) with a manual and two automated tests on Code 1. The MUT comes from an open-source project, Airsonic [21]. It has three string-type parameters, which are the user’s name, password, and email, respectively. The manual test assigns the values "Simon", "secret", and "simon@example.org" to them, matching the contexts of the three parameters. The input values generated by an LLM-based test generation tool (ChatUniTest [1]) are "testUser", "testPassword", and "test@example.com". They also match the contexts. Conversely, the input values from EvoSuite [2] are "|x45e*3q4+", " [stream]", and "". These values make the automated test less readable than the manual one. In our survey (see Sec. 5.3), $22$ out of $30$ respondents prefer the manual test to EvoSuite’s test regarding readability. The corresponding values from Randoop [3] are "hi!", "", and "", posing the same issue. Such deficiency prevails in tests generated by traditional tools. Fig. 1 aggregates the inputs from manual tests and tests generated by various testing tools for 30 classes (see Sec. 5.2). It shows that manual tests contain diversified real-world objects (e.g., person names, emails, and locations), while EvoSuite and Randoop [3] generate a large number of random inputs. It could be worse for methods with multiple parameters. One interviewee gave us the following feedback:
@@
The secondary objective complementing branch objectives proposed by Afshan et al. [10] guides genetic algorithms [5, 25] to generate inputs that are more likely to occur in the real language. However, this objective ignores the source code, thus ignoring the different input readability requirements of username and email. Hence, the optimized inputs may also not satisfy these readability requirements hidden in the source code. The approach of Deljouyi et al. [22] to extract readable test inputs from manual-written E2E tests can not applied in this JAVA class under test without extra manual efforts since the Airsonic project does not have E2E tests.
@@ 1.2 Contributions
To solve the challenges shown in the motivating example, this paper proposes a novel concept called readability context, which refers to the requirements for test readability described in the source code (e.g., valid person names for the username parameter in Code 1). Based on the readability context, this paper proposes the Context Consistency Criterion (C3). First, C3 passes detailed information (including the source code and the comments) to a Large Language Model (LLM) and requests it to mine the readability contexts for the primitive-type (including string-type) parameters. Besides, we also pass few-shots [27] (i.e., pre-defined similar questions and their answers for LLM to quickly adapt to this mine-context task) to LLM. Second, C3 checks how well the inputs of the tests are consistent with readability contexts based on LLMs, NLP tools, and regex matching.
@@
$\bullet$ We compare manual tests, two traditional test generation tools (EvoSuite [2] and Randoop [3]), EvoSuiteC3, and an LLM-based tool (ChatUniTest [1]) on $30$ JAVA classes with $115$ parameters that have readability contexts. Fig. 1 summarizes test inputs for these parameters and reveals that EvoSuiteC3 and ChatUniTest’s tests perform closely to manual ones in input readability. Meanwhile, EvoSuite and Randoop largely fail to generate readable string-type inputs (Sec. 5.2). Furthermore, we invite $30$ programmers with varied backgrounds to quantify the readability of manual and automated tests. The results show that programmers reach a consensus with C3 (Sec. 5.3).
The rest of this paper is organized as follows. Sec. 2 discusses the related work. Sec. 3 and 4 describe C3 and EvoSuiteC3’s system design, respectively. Sec. 5 illustrates our evaluation. Sec. 6 presents the threats to validity. Sec. 7 concludes the whole paper. Sec. 8 provides the data availability.
@@ 2 Related Work
Traditional test generation. This paper uses two traditional test generation tools, Randoop [3] and EvoSuite [2], to represent two methods. The first, random testing [3, 28, 29, 30], generates random test inputs to satisfy coverage criteria or expose faults in the software under test. Its lightweight nature allows for easy testing of large-scale systems, providing a solution to the scalability issues of static methods like model checking and symbolic execution [31, 32]. However, random testing has limitations. While it can cover source lines, the random inputs may lack readability. Our evaluation (Sec. 5) reveals that most string-type inputs generated by Randoop are meaningless (e.g., "hi!"), which can confuse developers. Additionally, random testing without explicit guidance is inefficient in generating high-coverage tests. The second method, search-based testing [33, 25, 34, 5, 2, 35], addresses this by using coverage criteria as fitness functions to guide genetic algorithms in generating test inputs. This approach may produce readable inputs if the optimization of a fitness function requires it (e.g., name == "Peter" in a condition statement). However, most readability requirements are not explicitly defined. Consequently, our experiments indicate that EvoSuite’s readability is close to Randoop’s.
@@
Readability of traditional test generation. Existing studies on the readability of traditional test generation can be divided into two categories. The first focuses on generating test summaries. Daka et al. [12] proposed summarizing a test’s covered goals as its name. Panichella et al. [13] suggested a more thorough summary incorporating covered goals, the JAVA class being tested, and additional details to create comments. Roy et al. [14] employed deep learning to name tests and internal variables and create comments. Li et al. [15] proposed an approach to generate test documentation. These summary-based approaches ignore the concrete inputs, which are also regarded as necessary for readability (see Sec. 5.3). The second category applies code readability studies to test readability. Based on Buse and Weimer’s research on code readability [36], Daka et al. [17] gathered a dataset of tests along with their human-rated readability, used common code metrics as features (including identifier length, parentheses, test length, and casts), and added extra features for tests (e.g., assertions). They trained a linear regression model to improve the readability of automated unit tests. Grano et al. [37] leveraged the existing code readability model [38] to compare the readability of manual and automated tests. Gregory Gay [16] proposed improving tests’ readability by leveraging LLMs to refactor tests, such as dividing a test into small ones and formatting test code to satisfy common code style standards. This paper closely aligns with the studies of Afshan et al. [10] and Deljouyi et al. [22]. Afshan et al. [10] incorporated a statistical language model into search-based test generation to predict the likelihood of the test input occurring in the real world and guide test generation as a secondary fitness function. Deljouyi et al. [22] proposed to leverage end-to-end (E2E) tests to generate understandable unit test inputs since an E2E test simulates a real user to execute an application’s entire workflow and may contain understandable inputs. However, this approach requires human effort to write E2E tests. All of the above-mentioned studies ignore the source code, which plays a crucial role in determining the readability of the test code. C3 complements these studies. It mines the readability contexts of the primitive-type parameter from the source code and judges whether a test input is consistent with the context as a readability indicator. Our experiments show that C3 identifies the shortcomings of EvoSuite and Randoop in readability, and EvoSuiteC3 improves EvoSuite’s readability.
@@
LLM-based test generation. Large language models (e.g., the GPT series [27]) have demonstrated capabilities across diverse domains, including unit test generation. Tang et al. [40] compared test suites generated by both LLM and EvoSuite in correctness, code coverage, complexity metrics [41], and bug detection. Lemieux et al. [42] suggested using LLMs to overcome plateaus in fitness functions during search-based test generation. AthenaTest [43] employs a pre-trained BART model [44] for test generation, while ChatUniTest [1] is an automated tool that guides LLMs in generating and repairing tests. More studies can be found in the survey paper [45]. Given its ready-to-use nature and ability to continuously guide LLM in generating syntactically correct tests, this paper uses ChatUniTest as a baseline for comparison with EvoSuite and Randoop. Many LLM studies [43, 1, 46] assert that tests generated by LLMs are more readable than those produced by traditional tools, a conclusion reached through manual review or developer interviews. Our approach, C3, offers a quantitative metric for measuring test readability automatically. Our experimental results validate their findings that LLMs surpass EvoSuite and Randoop in generating readable tests. We also find that LLMs outperform developers as their tests are readable and achieve higher code coverage than manual tests.
@@ 3 Context Consistency Criterion (C3)
This section describes the system design of Context Consistency Criterion (C3). Fig. 2 shows its overview. Without losing any generality, we use JAVA as the target language to introduce C3, while C3 is not tied to any specific language. C3 takes the tested source code (e.g., a JAVA class or a method) and the test code as the input. C3’s framework comprises two components, i.e., the Mining Component and the Judging Component. The Mining Component mines the readability contexts from the tested source code (Sec. 3.1). C3 extracts the primitive-type parameters from the source code. Then, for each parameter, C3 passes its text information (i.e., the source code and comments) to a Large Language Model (LLM) and lets LLM choose a context from a prepared context collection (shown in Table I). The Judging Component measures the consistency of test inputs with the readability contexts (Sec. 3.2). C3 extracts the primitive-type inputs from the test. Then, it uses various tools (i.e., LLM, NLP tools, and regex matching) to check whether an input satisfies its context as the readability result.
@@ 3.1 Mining Component
The first step is to mine the contexts from the code under test. C3 extracts the primitive-type parameters; then, C3 uses each parameter’s text information and the prepared contexts to ask an LLM which context it matches.
@@
Primitive-type parameter extractor. We leverage an open-source JAVA parser [47] to analyze the code and extract its parameters of primitive types, including string, byte, short, int, long, float, and double. Note that (1) the primitive types in this paper include string although JAVA regards it as an object type since it is common in nearly all real-world projects and can be initialized by literal values like other primitive types, and (2) we omit char and boolean since their range of values is small and finite, posing no issues with readability. Besides, for a parameter, we extract its method’s source and comments used to construct LLM’s prompt. Note that we only use the method source code instead of the whole class code, thus significantly reducing the prompt size and decreasing the model’s financial costs.
@@
Prepared contexts. We divide the concrete primitive types into two abstract types, i.e., String (including string) and Number (others), and introduce their contexts, respectively.
@@
(1) String. We use the named entity (a concept from NLP) [48, 49, 39, 50] to prepare contexts for String. Named entity refers to real-world context-aware objects, such as person names and emails. Referencing previous studies [39, 50], we prepare $23$ contexts shown in Table I, which can be grouped into six context categories: Cyberspace ($7$), Geographic and Abstract Location ($6$), Temporal ($4$), Number (not Number type, but Number as String contexts) ($3$), Personal Identifier ($2$), and Finance ($1$). Note that “LOCATION” and “TIME” are two generic contexts. We use them to capture potential contexts in the source code that belong to Geo-Abstract Location and Temporal but do not match any other contexts in these two categories. For example, “restaurant” does not match any specific contexts (e.g., “CITY” and “GPE”) of Geo-Abstract Location but can be captured by “LOCATION”.
@@
(2) Number. While named entities serve as String’s contexts, we employ literal formats for Number’s contexts, including base formats (e.g., 0b100, 0xfe) and Scientific Notation (e.g., 1.2e20f). We add two more contexts, “FIXEDLENGTH” and “LONGNUMBER”. The former targets numbers of fixed lengths, such as credit card numbers, while the latter is designed for potentially overly long numbers. Code 3 from our survey questions (Sec. 5.3) provides an example.
@@
We start illustrating the prompt-constructing process by introducing two parameters for controlling the prompt size. In LLM, a token is a basic unit of text [27], and an LLM has a token limit for a conversation. We use MaxToken to represent the token size limit. Then, we use Remain to keep the token space for the target question (i.e., the third part of Fig. 3) and LLM’s response. First, we calculate the RestToken by $(\textit{MaxToken}-(\textit{Remain}+strToken(\textit{targetQuestion})))$, where strToken computes the number of tokens within a string. Then, we add the role setting message to Messages and update RestToken. Next, we iterate to add a shot and update RestToken if there is enough token space. Finally, we add the targeting question. For the Number type, we have an extra operation. We observe that an important factor determining a Number parameter’s context is the operators it uses. The or method on Code 2 presents an example. The | operator denotes that the values of a and b should be in binary format, i.e., their contexts are “BINARY”. Therefore, we emphasize this factor by appending a sentence to the end of the targeting question, i.e., The operators involved by this parameter are [op1, ...]. Meanwhile, two tests identical for compilers, test1 and test2, vary in readability: The second is more readable since its inputs match the contexts.
@@
After generating messages, we pass them to LLM to get the result, which could be a concrete context or MISC. MISC means that LLM regards that no context matches this parameter.

### v2

@@ Abstract
Automated test generation tools (e.g., EvoSuite and Randoop) usually generate unit tests with higher code coverage than manual tests. However, the readability of automated tests is crucial for code comprehension and maintenance. The readability of unit tests involves many aspects. In this paper, we focus on test inputs. The central limitation of existing studies on input readability is that they focus on test codes alone without taking the tested source codes into consideration, making them either ignore different source codes’ different readability requirements or require manual efforts to write readable inputs. However, we observe that the source codes specify the contexts that test inputs must satisfy. Based on such observation, we introduce the Context Consistency Criterion (a.k.a, C3), which is a readability measurement tool that leverages Large Language Models to extract primitive-type (including string-type) parameters’ readability contexts from the source codes and checks whether test inputs are consistent with those contexts. We have also proposed EvoSuiteC3. It leverages C3’s extracted contexts to help EvoSuite generate readable test inputs. We have evaluated C3’s performance on $409$ JAVA classes and compared manual and automated tests’ readability under C3 measurement. The results are two-fold. First, The Precision, Recall, and F1-Score of C3’s mined readability contexts are $84.4\%$, $83\%$, and $83.7\%$, respectively. Second, under C3’s measurement, the string-type input readability scores of EvoSuiteC3, ChatUniTest (an LLM-based test generation tool), manual tests, and two traditional tools (EvoSuite and Randoop) are $90\%$, $83\%$, $68\%$, $8\%$, and $8\%$, showing the traditional tools’ inability in generating readable string-type inputs. We have conducted a survey based on the questionnaires collected from 30 programmers with varied backgrounds. The results reveal that when C3 identifies readable differences between tests, programmers tend to give similar opinions of the test’s readability of C3.
@@ 1 Introduction
Figure 1: Overview of test inputs from manual and automated tests for $115$ parameters from $30$ JAVA classes (see Sec.  5.2): Manual, EvoSuiteC3 (see Sec.  1), and ChatUniTest [1] tests contain many real-world entities, such as person names, emails, URLs, and locations; EvoSuite [2] and Randoop [3] tests contain many random inputs, as well as JAVA class names, JSON strings, and UUIDs [4] since these two tools tend to use predefined seeds and static or dynamic information of classes under test as test inputs
Unit testing is a popular testing technique among developers. A unit test checks whether a method inside the source code behaves functionally correctly. Existing unit test generation tools (e.g., search-based tool EvoSuite  [2, 5] and feedback-directed random tool Randoop [3]) generate unit tests that are close to and even outperform the manually written ones in many coverage criteria [6], such as branch coverage.
@@ 16 User user7 = new User("hi!", "", "");
Code 1: A constructor from Airsonic [12] with five tests differing in inputs
The readability of unit tests involves many aspects, including but not limited to tests’ method names [13], comments [14, 15], documents [16], identifiers, and assertions [17, 18]. In this paper, we focus on test inputs. Many studies [10, 19, 20, 21] and our survey (see Sec. 5.4) show that test inputs are crucial in readability. However, the random inputs generated by existing tools (EvoSuite and Randoop) cost understanding time. A former survey [10] showed that compared to strings occurring in natural languages (e.g., "mail"), random strings (e.g., "Qu5ua") require people at most $1.6$ times more comprehension time. To deal with this issue, Afshan et al. [10] used a statistical model to predict the likelihood of the test input occurring in the real world. They formulated the likelihood as a secondary objective to guide test generation. The main issue with this approach is that its model ignores the source code when judging the input readability. However, different source codes have different requirements for input readability. For example, a parameter email calls for email addresses as readable inputs, while city requires city names. Deljouyi et al. [22] proposed to extract understandable inputs from end-to-end (E2E) tests’ execution trace logs since an E2E test aims to simulate a real user to execute an application’s entire workflow [23, 24] and may contain understandable inputs. However, this approach requires human effort to write E2E tests.
@@ 1.1 A Motivating Example
We leverage an example to demonstrate the lack of readability of automated tests and the challenges of existing studies. This example is a simplified constructor method under test (MUT) with a manual and two automated tests on Code 1. The MUT comes from an open-source project, Airsonic [12]. It has three parameters, which represent the user’s name, password, and email, respectively. The manual test assigns the values "Simon", "secret", and "simon@example.org" to them, matching the contexts of the three parameters. The input values generated by an LLM-based test generation tool (ChatUniTest [1]) are "testUser", "testPassword", and "test@example.com". They also match the contexts. Conversely, the input values from EvoSuite [2] are "|x45e*3q4+", " [stream]", and "". These values make the automated test less readable than the manual one. In our survey (see Sec. 5.4), $22$ out of $30$ respondents prefer the manual test to EvoSuite’s test regarding readability. The corresponding values from Randoop [3] are "hi!", "", and "", posing the same issue. Such deficiency prevails in tests generated by traditional tools. Fig. 1 aggregates the inputs from manual tests and tests generated by various testing tools for 30 classes (see Sec. 5.2). It shows that manual tests contain diversified real-world objects (e.g., person names, emails, and locations), while EvoSuite and Randoop [3] generate a large number of random inputs. It could be worse for methods with multiple parameters. One interviewee gave us the following feedback:
@@
The secondary objective complementing branch objectives proposed by Afshan et al. [10] guides genetic algorithms [5, 25] to generate inputs that are more likely to occur in the real language. However, this objective ignores the source code, thus ignoring the different input readability requirements of username and email. Hence, the optimized inputs may also not satisfy these readability requirements hidden in the source code. The approach of Deljouyi et al. [22] to extract readable test inputs from manual-written E2E tests could not applied in this JAVA class under test without extra manual efforts since the Airsonic project does not have E2E tests.
@@ 1.2 Contributions
To solve the challenges shown in the motivating example, this paper proposes a novel concept called readability context, which refers to the requirements for test readability described in the source code (e.g., valid person names for the username parameter in Code 1). Based on the readability context, this paper proposes the Context Consistency Criterion (C3). First, C3 passes detailed information (including the source code and the comments) to a Large Language Model (LLM) and requests it to mine the readability contexts for the primitive-type parameters. We skip those complex-type parameters (e.g., object-types excluding string-type) since we only focus on the readability of literal input values (e.g., number and string literals) in this paper. Besides, we also pass few-shots [27] (i.e., pre-defined similar questions and their answers for LLM to quickly adapt to this mine-context task) to LLM. Second, C3 checks how well the inputs of the tests are consistent with readability contexts based on LLMs, NLP tools, and regex matching.
@@
$\bullet$ We compare manual tests, two traditional test generation tools (EvoSuite [2] and Randoop [3]), EvoSuiteC3, and an LLM-based tool (ChatUniTest [1]) on $30$ JAVA classes with $115$ parameters that have readability contexts. Fig. 1 summarizes test inputs for these parameters and reveals that EvoSuiteC3 and ChatUniTest’s tests perform closely to manual ones in input readability. Meanwhile, EvoSuite and Randoop largely fail to generate readable string-type inputs (Sec. 5.2 and Sec. 5.3). Furthermore, we invite $30$ programmers with varied backgrounds to quantify the readability of manual and automated tests. The results show that programmers reach a consensus with C3 (Sec. 5.4).
The rest of this paper is organized as follows. Sec. 2 discusses the related work. Sec. 3 and 4 describe C3 and EvoSuiteC3’s system design, respectively. Sec. 5 illustrates our evaluation. Sec. 6 shows how to extend C3. Sec. 7 presents the threats to validity. Sec. 8 concludes the whole paper. Sec. 9 provides the data availability.
@@ 2 Related Work
Traditional test generation. This paper uses two traditional test generation tools, Randoop [3] and EvoSuite [2], to represent two methods. The first, random testing [3, 28, 29, 30], generates random test inputs to satisfy coverage criteria or expose faults in the software under test. Its lightweight nature allows for easy testing of large-scale systems, providing a solution to the scalability issues of static methods like model checking and symbolic execution [31, 32]. However, random testing has limitations. While its tests cover source lines, the random inputs may lack readability. Our evaluation (Sec. 5) reveals that most string-type inputs generated by Randoop are meaningless (e.g., "hi!"), which confuse developers. Additionally, random testing without explicit guidance is inefficient in generating high-coverage tests. The second method, search-based testing [33, 25, 34, 5, 2, 35], addresses this by using coverage criteria (e.g., branch coverage) as fitness functions to guide genetic algorithms in generating test inputs. This approach may produce readable inputs if covering a criterion goal requires a readable input. For example, If a test covers the true branch of the branch condition userName == "Peter", it contains a readable input ("Peter") for a person’s name. However, most readability requirements are not explicitly defined in real-world projects. Consequently, our experiments indicate that EvoSuite’s readability is close to Randoop’s (Sec. 5.2).
@@
Readability of traditional test generation. We divide existing studies on the readability of traditional test generation into two categories. The first focuses on generating test summaries. Daka et al. [13] proposed summarizing a test’s covered goals as its name. Panichella et al. [14] suggested a more thorough summary incorporating covered goals, the JAVA class being tested, and additional details to create comments. Roy et al. [15] employed deep learning to name tests and internal variables and create comments. Li et al. [16] proposed an approach to generate test documentation. These summary-based approaches ignore the concrete inputs, which are also regarded as necessary for readability (see Sec. 5.4). The second category applies code readability studies to test readability. Based on Buse and Weimer’s research on code readability [36], Daka et al. [18] gathered a dataset of tests along with their human-rated readability, used common code metrics as features (including identifier length, parentheses, test length, and casts), and added extra features for tests (e.g., assertions). They trained a linear regression model to improve the readability of automated unit tests. Grano et al. [37] leveraged the existing code readability model [38] to compare the readability of manual and automated tests. Gregory Gay [17] proposed improving tests’ readability by leveraging LLMs to refactor tests, such as dividing a test into small ones and formatting test code to satisfy common code style standards. This paper closely aligns with the studies of Afshan et al. [10] and Deljouyi et al. [22]. Afshan et al. [10] incorporated a statistical language model into search-based test generation to predict the likelihood of the test input occurring in the real world and guide test generation as a secondary fitness function. Deljouyi et al. [22] proposed to leverage end-to-end (E2E) tests to generate understandable unit test inputs since an E2E test simulates a real user to execute an application’s entire workflow and may contain understandable inputs. However, this approach requires human effort to write E2E tests. All of the above-mentioned studies ignore the source code, which plays a crucial role in determining the readability of the test code. C3 complements these studies. It mines the readability contexts of the primitive-type parameter from the source code and judges whether a test input is consistent with the context as a readability indicator. Our experiments show that C3 identifies the shortcomings of EvoSuite and Randoop in readability, and EvoSuiteC3 improves EvoSuite’s readability.
@@
LLM-based test generation. Large language models (e.g., the GPT series [27]) have demonstrated capabilities across diverse domains, including unit test generation. Tang et al. [40] compared test suites generated by both LLM and EvoSuite in correctness, code coverage, complexity metrics [41], and bug detection. Lemieux et al. [42] suggested using LLMs to overcome plateaus in fitness functions during search-based test generation. Liu et al. [43] proposed to leverage LLMs to improve GUI testing’s page coverage. AthenaTest [44] employs a pre-trained BART model [45] for test generation, while ChatUniTest [1] is an automated tool that guides LLMs in generating and repairing tests. More studies are in the survey paper [46]. Given its ready-to-use nature and ability to continuously guide LLM in generating syntactically correct tests, this paper uses ChatUniTest as a baseline for comparison with EvoSuite and Randoop. Many LLM studies [44, 1, 47] assert that tests generated by LLMs are more readable than those produced by traditional tools, a conclusion reached through manual review or developer interviews. Our approach, C3, offers a quantitative metric for measuring test readability automatically. Our experimental results validate their findings that LLMs surpass EvoSuite and Randoop in generating readable tests. We also find that LLMs outperform developers as their tests are readable and achieve higher code coverage than manual tests.
@@ 3 Context Consistency Criterion (C3)
This section describes the design details of Context Consistency Criterion (C3). Fig. 2 shows its overview. Without losing any generality, we use JAVA as the target language to introduce C3, while C3 is not tied to any specific language. C3 takes the tested source code (e.g., a JAVA class or a method) and the test code as the input. C3’s framework comprises two components, i.e., the Mining Component and the Judging Component. The Mining Component mines the readability contexts from the tested source code (Sec. 3.1). C3 extracts the primitive-type parameters from the source code. Then, for each parameter, C3 passes its text information (i.e., the source code and comments) to a Large Language Model (LLM) and lets the LLM choose a context from a group of prepared contexts (shown in Table I). The Judging Component measures the consistency of test inputs with the readability contexts (Sec. 3.2). C3 extracts the primitive-type inputs from the test. Then, it uses various tools (i.e., LLM, NLP tools, and regex matching) to check whether an input satisfies its context as the readability result.
@@ 3.1 Mining Component
The first step is to mine the contexts for each primitive-type parameter from the code under test. A simple method is to directly pass each parameter’s text information (the source code and comments) to LLM and ask, What context should this parameter’s value match? However, such a method has two major flaws. First, an LLM may hallucinate [48] when facing an open-ended question, i.e., it may provide a fictitious response. For example, VHTest [49] (an LLM benchmark) showed that, on average, $925$ of $1200$ open-ended questions induce four LLMs to hallucinate. Second, an LLM will return any possible result for such a question. Even if we regard this unpredictable result as correct, it is difficult to leverage it to judge tests’ readability (Sec. 3.2) and improve existing tools (Sec. 1). Hence, we prepare a group of contexts (Table I) in advance so that we can restrict LLM to choosing only one of these contexts or answering no context. Our experiment (Sec. 3.1) showed that in a sample of $4485$ parameters, $1048$ ($23.4\%$) parameters were confirmed to be with our prepared contexts. Our design is extensible if end users add more specific contexts, and we provide a guide instruction (Sec. 6) for users to add more contexts.
@@
Primitive-type parameter extractor. We leverage an open-source JAVA parser [50] to analyze the code and extract its parameters of primitive types, including string, byte, short, int, long, float, and double. Although JAVA regards string as an object type, the primitive types in this paper include it because string can be initialized by literal values like other primitive types. We omit char and boolean since the possible values of char are 00 to $65536$ and boolean’s value is only either true or false. Hence, their range of concrete values is finite and small, posing no issues with readability. Besides, for a parameter, we extract its method’s source and comments used to construct LLM’s prompt. Note that we only use the method source code instead of the whole class code, thus significantly reducing the prompt size and decreasing the model’s financial costs.
@@
Prepared contexts. We divide the concrete primitive types into two groups, i.e., String (including string) and Number (byte, short, int, long, float, and double). Then, we introduce their contexts by group.
@@
(1) String. Named Entity Recognition (NER) [51, 52, 39, 53] is a natural language processing task that identifies key information in the text and classifies it into predefined categories (called named entities). A named entity refers to a real-world context-aware object, such as a person’s name, organization, and email. We adopt named entities as the contexts for String. Referencing previous studies [39, 53], we prepare $23$ contexts shown in Table I, which are grouped into six context categories: Cyberspace ($7$), Geographic and Abstract Location ($6$), Temporal ($4$), Number (not Number type, but Number as String contexts) ($3$), Personal Identifier ($2$), and Finance ($1$). Note that “LOCATION” and “TIME” are two generic contexts. We use them to capture potential contexts in the source code that belong to Geo-Abstract Location and Temporal but do not match any other contexts in these two categories. For example, “restaurant” does not match any specific contexts (e.g., “CITY” and “GPE”) of Geo-Abstract Location but is captured by “LOCATION”.
@@
(2) Number. While named entities serve as String’s contexts, we employ literal formats for Number’s contexts, including base formats (e.g., 0b100, 0xfe) and Scientific Notation (e.g., 1.2e20f). We add two more contexts, “FIXEDLENGTH” and “LONGNUMBER”. The former targets numbers of fixed lengths, such as credit card numbers, while the latter is designed for potentially overly long numbers.
It is worth noting that users can add extra contexts to meet their needs. Sec. 6 shows how to add new contexts.
@@
We start illustrating the prompt-constructing process by introducing two parameters for controlling the prompt size. In LLM, a token is a basic unit of text [27], and an LLM has a token limit for a conversation. We use MaxToken to represent the token size limit. Then, we use Remain to keep the token space for LLM’s response. This prompt-constructing process aims to add shots as many as possible. First, we calculate the RestToken (to record the remaining token space) by $(\textit{MaxToken}-(\textit{Remain}+strToken(\textit{targetQuestion})))$, where strToken computes the number of tokens within a string. Then, we add the role setting message to Messages and update RestToken. Next, we iterate to add a shot and update RestToken if there is enough token space. Finally, we add the targeting question. For the Number type, we have an extra operation. We observe that an important factor determining a Number parameter’s context is the operators it uses. The or method on Code 2 presents an example. The | operator denotes that the values of a and b should be in binary format, i.e., their contexts are “BINARY”. Therefore, we emphasize this factor by appending a sentence to the end of the targeting question, i.e., The operators involved by this parameter are [op1, ...]. Meanwhile, two tests identical for compilers, test1 and test2, vary in readability: The second is more readable since its inputs match the contexts. After generating messages, we pass them to LLM to get the result, which could be a concrete context or MISC. MISC means that LLM regards that no context matches this parameter.